# Better wit than wealth: Dynamic Parametric Retrieval Augmented Generation for Test-time Knowledge Enhancement

### v1

@@ 1 Introduction
Knowledge-intensive tasks, such as question answering (QA), require leveraging extensive world and domain knowledge [18; 20]. While Large Language Models (LLMs) have demonstrated impressive capabilities across various tasks and fields [25], they struggle to handle knowledge-intensive tasks independently due to limitations in accessing external knowledge and a tendency to generate hallucinated information [8].
To mitigate incorrect internal knowledge and improve performance in QA, researchers have developed knowledge-augmented methods for LLMs. One of the most used methods called Retrieval-Augmented Generation (RAG) [11], which retrieves relevant documents from external sources (e.g., Wikipedia) and integrates both the retrieved documents and the query into the model’s input [15] (as shown in Figure 1 (a)).
Standard RAG methods [3; 16; 34] share a common approach: injecting external knowledge into LLMs by appending retrieved documents to the input context, referred to as in-context injection. This technique harnesses the reasoning capabilities of LLMs to address knowledge gaps and reduce misinformation [5]. However, it introduces several challenges. As the number and length of retrieved documents increase, the extended context leads to significantly higher inference costs. Additionally, when the internal knowledge of LLMs conflicts with the external context, a phenomenon known as RAG hallucination (or knowledge conflict) arises, causing LLMs to produce incorrect answers even when accurate documents are provided [48; 37].
Parametric RAG (PRAG) employs another way of injecting knowledge which integrates external knowledge directly into the parameters of LLMs, known as parameter injection (in Figure 1 (b)). PRAG first rewrites retrieved documents and transforms them into QA pairs, facilitating knowledge memorization and manipulation [1; 2]. The workflow of PRAG is divided into two stages. During the offline phase, these augmented documents are fine-tuned with LoRA [14], encoding contextual knowledge directly into parameters. In the online phase, retrieved documents are replaced with loadable parameters, reducing context length while embedding external knowledge into the internal representation of LLMs for test-time knowledge enhancement. However, this approach has significant trade-offs. Augmenting, training, and storing parameters for each retrieved document incurs high computational and storage costs, coupled with non-generalization, severely limiting scalability in real-world applications (e.g., systems in domains requiring frequent knowledge updates or multi open-domain QA tasks). This presents a critical challenge for PRAG: how can we minimize the overhead of converting documents into parametric knowledge without compromising performance?
@@
In this work, we propose Dynamic Parametric RAG (DyPRAG), a lightweight framework that converts documents into parameters directly, making DyPRAG seamlessly enhance the knowledge of LLMs in a plug-and-play manner at test-time (as shown in Figure 1 (c)). Rethinking of PRAG, its intrinsic goal is to obtain an underlying mapping function $\mathcal{F}$ which transforms external documents into parameters through augmentation and training for each document. We believe better wit than wealth—the ability to convert documents into parameters holds greater significance than arduously storing a large collection of parametric representations. Building on this insight, we aim to establish a more generalized mapping function by directly modeling $\mathcal{F}$ with a small hypernetwork $\mathcal{F}^{\prime}_{\phi}$, referred to as parameter translator. With well-trained $\mathcal{F}^{\prime}_{\phi}$, DyPRAG eliminates the need for the cumbersome training and storage process required by PRAG. After a detailed analysis of computation and storage overhead, our method significantly reduces the high inference costs of traditional RAG while eliminating the rigid training and storage costs of PRAG.
@@
Through extensive experiments, we derive two key findings: 1) DyPRAG enhances the test-time parametric knowledge of LLMs effectively. During evaluation, DyPRAG outperforms standard RAG across different scales of LLMs, demonstrating its ability to enhance the internal knowledge of LLMs. Furthermore, although PRAG learns $\mathcal{F}$ by training separately for each document, DyPRAG achieves comparable or even better performance in various scenarios with significantly lower costs. 2) Combining test-time generated parametric knowledge with contextual knowledge leads to superior knowledge fusion. Following  [36], we further investigate the combination of in-context injection and parameter injection, referred to as DyPRAG-Combine. In both independent identically distributed and out-of-distribution settings, DyPRAG-Combine achieves the best results, showing strong generalization ability. Additionally, we find that DyPRAG-Combine effectively relieves RAG hallucination by first injecting context-related parameters to reduce the knowledge conflicts. Based on the experimental results on RAGTruth [45] benchmark, we observe that DyPRAG-Combine enables LLMs to internalize previous unseen knowledge. This suggests that integrating parametric and contextual knowledge using DyPRAG could be a powerful approach for building a trustworthy RAG system in real-world applications.
In conclusion, the contributions of this paper are summarized as follows:
@@ 2.1 Retrieval Augmented Generation
Large language models (LLMs) have demonstrated remarkable performance across diverse applications. However, their inherent knowledge often falls short in handling knowledge-intensive tasks, highlighting the need for external knowledge integration to ensure robust performance in such contexts. A prominent approach to bridging this gap is Retrieval-Augmented Generation (RAG), which enhances LLMs by incorporating relevant external knowledge sources [4; 41; 42; 12]. The retrieved documents are appended to the LLM’s input context, enabling it to leverage knowledge beyond its training data [22]. We refer to this as the in-context injection paradigm. However, this approach leads to high inference costs and can exceed context length limits as the number and length of retrieved documents increase [46]. To address this issue, recent study introduces Parametric RAG (PRAG) [36], a paradigm that fine-tunes the model on augmented documents, encoding useful information into its parameters. While PRAG mitigates the inference cost, it introduces additional training and storage costs due to the need to obtain and store LoRA parameters. Our proposed method significantly reduces the high costs associated with standard RAG and PRAG while achieving superior generalization. By combining contextual knowledge with test-time generated parametric knowledge through DyPRAG, it enables LLMs to better manipulate knowledge, effectively mitigating RAG hallucination. This makes it a more practical and robust solution for real-world applications.
@@ 2.2 Context Compression
Context compression is widely adopted to improve the efficiency of LLMs in processing contextual knowledge. Recent studies [30; 9] propose condensing long contexts into soft prompts, allowing LLMs to utilize information more effectively. Meanwhile, other works  [27; 43] focus on transforming context chunk into LoRA modules to improve the understanding ability of extended contexts while xRAG [6] integrates context compression by mapping documents into a compact token representation. Similarly, AAG [24] draws inspiration from human cognition, retrieving and recalling relevant knowledge to compensate for knowledge gaps. This approach activates relevant information within LLMs without relying on external resources. Our proposed method dynamically converts retrieved documents into parameters, effectively unifying internal and external knowledge while mitigating RAG hallucination issue. Furthermore, it functions as a plug-and-play solution to enhance the efficiency and adaptability of standard RAG systems by test-time parametric knowledge enhancement.
@@ 3 Methodology
Figure 2: An illustration of the DyPRAG method. In the offline phase, Stage 1 follows the same parameterization process as PRAG to collect Doc-Param pairs. In Stage 2, we train the parameter projector $\mathcal{F}^{\prime}_{\phi}$ to learn the mapping function from documents to parameters. During the online Stage 3, the trained $\mathcal{F}^{\prime}_{\phi}$ dynamically generates LoRA modules, enabling LLMs to enhance their knowledge at test-time.
@@ 3.2 Dynamic Parametric RAG
where we calucate the overall alignment loss for each document $d_{i}$, $\lambda_{1}$ and $\lambda_{2}$ are tunable hyper-parameter which set to 100 and 0.01 separately to make loss range similar.
@@
Table 1: Comparison of cost metrics for different baselines. ATTN denotes the time complexity of the self-attention module as $O(|I|^{2}h)$, and FFN represents the FFN with $O(|I|h^{2})$, where context length $|I|=1$ and $|R|$ denotes the response length. ii indicates significant high cost, aa denotes negligible cost, and aa represents temporal storage.
Method Inference Cost Training Cost Storage Cost
RAG $|R|\times((c|d|+|q|)^{2}\times\text{ATTN}+(c|d|+|q|)\times\text{FFN})$
[…]
@@ 4 Experiments
DyPRAG (ours)  35.83  44.89  14.81  8.64  25.18  13.72  41.39  19.67  6.64  31.94  24.27
DyPRAG-Combine (ours)  43.03  47.20  17.04  8.55  26.46  21.56  41.25  27.57  22.69  33.57  28.89
LLaMA3-8B Standard RAG  58.43  47.77  19.20  11.07  34.55  19.68  42.10  24.23  16.13  35.45  30.86
@@
4.1 Experiments Details
Datasets.
We assess our approach using a variety of benchmark datasets, each specifically crafted to evaluate distinct reasoning abilities, including multi-hop reasoning and commonsense inference. The selected datasets are as follows:
@@ 4.1 Experiments Details
For evaluation, we use the F1 score (%), which balances precision and recall by accounting for partially correct answers. Both 2WQA and HQA categorize questions by reasoning type, with 2WQA having four categories and HQA two. To compare DyPRAG with other RAG baselines across reasoning tasks, Table 2 and 5 report performance on each sub-task separately, using the first 300 questions from each sub-dataset, following [36].
@@ 4.3 Main Results
In this section, we present the main experimental result and in-depth analysis of DyPRAG compared with standard RAG, efficient RAG, PRAG baselines and a combined setting that leverages both parametric and contextual knowledge.
@@
Since PRAG learns the mapping function $\mathcal{F}$ by training separately for each document, it can be considered as a performance upper bound of DyPRAG. Remarkably, our proposed DyPRAG achieves comparable or even superior results across various tasks as shown in Table 2. For instance, using LLaMA3.2-1B, DyPRAG achieves an average score of 27.57%, surpassing PRAG by 1.06% and standard RAG by 0.58%. A notable result is observed in the bridge sub-task of 2WQA, where DyPRAG achieves a score of 48.15%, outperforming RAG and PRAG by a significant margin of 21.37% and 23.81%, respectively. This demonstrates that our method exhibits robust generalization capabilities when trained on diverse datasets. We also compare our DyPRAG framework with several efficient RAG baselines, including FLARE [17] and DRAGIN [35]. As shown in Table 5, both DRAGIN and FLARE outperform standard RAG in most 2WQA settings. However, DyPRAG achieves even better results, demonstrating its superiority. For example, when using LLaMA3-8B as the base model, DyPRAG outperforms DRAGIN and FLARE by 1.56% and 2.63% on 2WQA, respectively. These results highlight the consistent performance improvements offered by DyPRAG over other efficient RAG methods, underscoring its effectiveness for test-time knowledge enhancement.
Performance deficiencies.
In experiments with Qwen2.5-1.5B (LLaMA3.2-8B), our DyPRAG averagely outperforms standard RAG by 1.10% (5.77%), but lags behind the upper-bound PRAG by 2.87% (4.77%). The performance gap primarily lies in the PQA task, where DyPRAG struggles to optimize due to the factual nature of the questions, which are predominantly occupation-related. The results indicate that single DyPRAG encounters challenges when processing excessively fine-grained knowledge. However, it still outperforms standard RAG significantly, demonstrating the efficiency of parameter injection at test-time.
@@
LLaMA3.2-1B Standard RAG 40.38 27.67
@@
Qwen2.5-1.5B Standard RAG 30.52 39.00
@@
LLaMA3-8B Standard RAG 43.27 45.67
@@
By combining in-context injection with parameter injection, DyPRAG-Combine achieves the best performance across all models, outperforming both efficient RAG and PRAG baselines. For instance, DyPRAG-Combine outperforms PRAG-Combine by 1.86% on LLaMA3.2-1B, 1.84% on Qwen2.5-1.5B and 1.08% on LLaMA3-8B on average. These results demonstrate that the dynamic parameters introduced by our approach effectively intergrade with contextual knowledge, enabling the two information sources to complement each other.
@@ 4.4 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalization ability of the DyPRAG method, we evaluate it in an out-of-distribution (OOD) scenario. Table 3 presents the OOD performance on the StrategyQA (SQA)[10] and IIRC[7] datasets utilizing first 300 test examples. Compared to the standard RAG model, our DyPRAG-Combine which incorporates document-related parametric knowledge can answer these questions more correctly, achieving the best performance across all scenarios. For example, DyPRAG-Combine improves performance on SQA by 22.66% using LLaMA3.2-1B and on IIRC by 13.63% using LLaMA3-8B.
These experimental results highlight the significant generalization ability of DyPRAG, surpassing that of PRAG, which is unable to handle this OOD scenario without additional offline training. Detailed implementation is provided in Appendix B.1.
4.5 Varying Number of Injected Documents
For standard RAG, the number of retrieved documents, denoted as $c$, is a crucial hyperparameter to tune. Recent studies [21; 44] have investigated the impact of longer context lengths on standard RAG. However, the effect of the number of injected documents in parametric form remains underexplored. Our proposed DyPRAG framework can seamlessly adapt to this scenario due to its inherent flexibility.

### v2

# Dynamic Parametric Retrieval Augmented Generation for Test-time Knowledge Enhancement

@@ 1 Introduction
Knowledge-intensive tasks, such as question answering (QA), require leveraging extensive world and domain knowledge [20; 22]. While Large Language Models (LLMs) have demonstrated impressive capabilities across various tasks and fields [27], they struggle to handle knowledge-intensive tasks independently due to limitations in accessing external knowledge and a tendency to generate hallucinated information [10].
To mitigate incorrect internal knowledge and improve performance in QA, researchers have developed knowledge-augmented methods for LLMs. One of the most used methods called Retrieval-Augmented Generation (RAG) [13], which retrieves relevant documents from external sources (e.g., Wikipedia) and integrates both the retrieved documents and the query into the model’s input [17] (as shown in Figure 1 (a)). Standard RAG methods [3; 18; 37] share a common approach: injecting external knowledge into LLMs by appending retrieved documents to the input context, referred to as in-context injection. This technique harnesses the reasoning capabilities of LLMs to address knowledge gaps and reduce misinformation [5]. However, it introduces several challenges. As the number and length of retrieved documents increase, the extended context leads to significantly higher inference costs. Additionally, when the internal knowledge of LLMs conflicts with the external context, a phenomenon known as RAG hallucination (or knowledge conflict) arises, causing LLMs to produce incorrect answers even when accurate documents are provided [51; 40].
@@
Parametric RAG (PRAG) employs another way of injecting knowledge which integrates external knowledge directly into the parameters of LLMs, known as parameter injection (in Figure 1 (b)). PRAG first rewrites retrieved documents and transforms them into QA pairs, facilitating knowledge memorization and manipulation [1; 2]. The workflow of PRAG is divided into two stages. During the offline phase, these augmented documents are fine-tuned with LoRA [16], encoding contextual knowledge directly into parameters. In the online phase, retrieved documents are replaced with loadable parameters. However, this approach has significant trade-offs. Augmenting, training, and storing parameters for each retrieved document incurs high computational and storage costs, coupled with non-generalization, severely limiting scalability in real-world applications (e.g., systems in domains requiring frequent knowledge updates or multi open-domain QA tasks). This presents a critical challenge for PRAG: how can we minimize the overhead of converting documents into parametric knowledge without compromising performance?
@@
In this work, we propose Dynamic Parametric RAG (DyPRAG), a lightweight framework that converts documents into parameters directly, making DyPRAG seamlessly enhance the knowledge of LLMs in a plug-and-play manner at test-time (as shown in Figure 1 (c)). Rethinking of PRAG, its intrinsic goal is to obtain an underlying mapping function $\mathcal{F}$ which transforms external documents into parameters through augmentation and training for each document. Building on this insight, we aim to establish a more generalized mapping function by directly modeling $\mathcal{F}$ with a small hypernetwork $\mathcal{F}^{\prime}_{\phi}$, referred to as parameter translator. With well-trained $\mathcal{F}^{\prime}_{\phi}$, DyPRAG eliminates the need for the cumbersome training and storage process required by PRAG. After a detailed analysis of computation and storage overhead, our method significantly reduces the high inference costs of traditional RAG while eliminating the rigid training and storage costs of PRAG.
@@
Through extensive experiments, we derive two key findings: 1) DyPRAG enhances the test-time parametric knowledge of LLMs effectively. During evaluation, DyPRAG outperforms standard RAG across different scales of LLMs, demonstrating its ability to enhance the internal knowledge of LLMs. Furthermore, although PRAG learns $\mathcal{F}$ by training separately for each document, DyPRAG achieves comparable or even better performance in various scenarios with significantly lower costs. 2) Combining test-time generated parametric knowledge with contextual knowledge leads to superior knowledge fusion. Following  [39], we further investigate the combination of in-context injection and parameter injection, referred to as DyPRAG-Combine. In both independent identically distributed and out-of-distribution settings, DyPRAG-Combine achieves the best results, showing strong generalization ability. Additionally, we find that DyPRAG-Combine effectively relieves RAG hallucination by first injecting context-related parameters to reduce the knowledge conflicts. Based on the experimental results on RAGTruth [33] benchmark, we observe that DyPRAG-Combine enables LLMs to internalize previous unseen knowledge. This suggests that integrating parametric and contextual knowledge using DyPRAG could be a powerful approach for building a trustworthy RAG system in real-world applications. We summarize our contributions as follows:
@@ 2.1 Retrieval Augmented Generation
Large language models (LLMs) have demonstrated remarkable performance across diverse applications. However, their inherent knowledge often falls short in handling knowledge-intensive tasks, highlighting the need for external knowledge integration to ensure robust performance in such contexts. A prominent approach to bridging this gap is Retrieval-Augmented Generation (RAG), which enhances LLMs by incorporating relevant external knowledge sources [4; 44; 45; 14]. The retrieved documents are appended to the LLM’s input context, enabling it to leverage knowledge beyond its training data [24]. We refer to this as the in-context injection paradigm. However, this approach leads to high inference costs and can exceed context length limits as the number and length of retrieved documents increase [48]. To address this issue, recent study introduces Parametric RAG (PRAG) [39], a paradigm that fine-tunes the model on augmented documents, encoding useful information into its parameters. While PRAG mitigates the inference cost, it introduces additional training and storage costs due to the need to obtain and store LoRA parameters. Our proposed method significantly reduces the high costs associated with standard RAG and PRAG while achieving superior generalization. By combining contextual knowledge with test-time generated parametric knowledge through DyPRAG, it enables LLMs to better manipulate knowledge, effectively mitigating RAG hallucination. This makes it a more practical and robust solution for real-world applications.
@@ 2.2 Context Compression
Context compression is widely adopted to improve the efficiency of LLMs in processing contextual knowledge. Recent studies [32; 11; 8] propose condensing long contexts into soft prompts, allowing LLMs to utilize information more effectively. Meanwhile, other works  [29; 46] focus on transforming context chunk into LoRA modules to improve the understanding ability of extended contexts while xRAG [7] integrates context compression by mapping documents into a compact token representation. Similarly, AAG [26] draws inspiration from human cognition, retrieving and recalling relevant knowledge to compensate for knowledge gaps. This approach activates relevant information within LLMs without relying on external resources. Our proposed method dynamically converts retrieved documents into parameters, effectively unifying internal and external knowledge while mitigating RAG hallucination issue. Furthermore, it functions as a plug-and-play solution to enhance the efficiency and adaptability of standard RAG systems by test-time parametric knowledge enhancement.
@@ 3 Methodology
Figure 2: An illustration of the DyPRAG method. In the offline phase, Stage 1 follows the same parameterization process as PRAG to collect Doc-Param pairs. In Stage 2, we train the parameter translator $\mathcal{F}^{\prime}_{\phi}$ to learn the mapping function from documents to parameters. During the online Stage 3, the trained $\mathcal{F}^{\prime}_{\phi}$ dynamically generates LoRA modules to enhance LLMs knowledge at test-time.
@@ 3.2 Dynamic Parametric RAG
where we calculate the overall alignment loss for each document $d_{i}$, $\lambda_{1}$ and $\lambda_{2}$ are tunable hyper-parameter which set to 100 and 0.01 separately to make loss range similar.
@@ 4 Experiments
4.1 Experiments Details
Datasets.
We assess our approach using a variety of benchmark datasets, each specifically crafted to evaluate distinct reasoning abilities, including multi-hop reasoning and commonsense inference. The selected datasets are 2WikiMultihopQA (2WQA) [15], HotpotQA (HQA) [50], PopQA (PQA) [28] and ComplexWebQuestions (CWQ) [41]. We provide detailed information about these datasets in Appendix B.1.
@@ 4.1 Experiments Details
DyPRAG (ours)  43.03  47.20  17.04  8.55  26.46  13.72  41.39  19.67  6.64  31.94  25.56
DyPRAG-Combine (ours)  35.83  44.89  14.81  8.64  25.18  21.56  41.25  27.57  22.69  33.57  27.60
LLaMA3-8B Vanilla  54.90  55.20  24.59  14.43  33.02  19.00  45.63  21.29  7.96  42.44  31.85
Standard RAG  58.43  47.77  19.20  11.07  34.55  19.68  42.10  24.23  16.13  35.45  30.86
@@
For evaluation, we use the F1 score (%), which balances precision and recall by accounting for partially correct answers. Both 2WQA and HQA categorize questions by reasoning type, with 2WQA having four categories and HQA two. To compare DyPRAG with other RAG baselines across reasoning tasks, Table 1 and 6 report performance on each sub-task separately, using the first 300 questions from each sub-dataset, following [39].
@@ 4.2 Baselines
  • Vanilla represents the answer from original LLMs without any external knowledge.
@@ 4.3 Main Results
In this section, we present the main experimental result and in-depth analysis of DyPRAG compared with vanilla model, standard RAG, efficient RAG, PRAG baselines and a combined setting . Notably, the vanilla model occasionally outperforms RAG in certain situations. We analyze the reasons for this in Appendix D, and confirm that it won’t affect the subsequent analysis.
@@
Since PRAG learns the mapping function $\mathcal{F}$ by training separately for each document, it can be considered as a performance upper bound of DyPRAG. Remarkably, our proposed DyPRAG achieves comparable or even superior results across various tasks as shown in Table 1. For instance, using LLaMA3.2-1B, DyPRAG achieves an average score of 27.57%, surpassing PRAG by 1.06%, standard RAG by 0.58% and vanilla model by 5.18%. A notable result is observed in the bridge sub-task of 2WQA, where DyPRAG achieves a score of 48.15%, outperforming RAG and PRAG by a significant margin of 21.37% and 23.81%, respectively. This demonstrates that our method learns more useful information when trained on diverse datasets. We also compare our DyPRAG framework with several efficient RAG baselines, including FLARE [19] and DRAGIN [38]. As shown in Table 6, both DRAGIN and FLARE outperform standard RAG in most 2WQA settings. However, DyPRAG achieves even better results, demonstrating its superiority. For example, when using LLaMA3-8B as the base model, DyPRAG outperforms DRAGIN and FLARE by 1.56% and 2.63% on 2WQA, respectively. These results highlight the consistent performance improvements offered by DyPRAG over other efficient RAG methods, underscoring its effectiveness for test-time knowledge enhancement.
@@
LLaMA3.2-1B Vanilla 10.99 21.67
Standard RAG 40.38 27.67
DyPRAG 14.04 39.67
@@
Qwen2.5-1.5B Vanilla 8.78 1.00
Standard RAG 30.52 39.00
DyPRAG 10.23 15.67
@@
LLaMA3-8B Vanilla 13.23 33.33
Standard RAG 43.27 45.67
DyPRAG 18.16 45.67
@@
Table 3: Ablation study of intermediate dimension $p$ of $\mathcal{F}^{\prime}_{\phi}$. The backbone model is the Qwen2.5-1.5B. The inference time is computed by average time of CWQ with batch_size of 1. The encode time is highlighted in red, while the translate time is marked in blue.
Method
CWQ
F1
Inference Time
Storage Cost
Vanilla 26.47 0.56 (0.47x) -
RAG 28.32 1.20 (1x) -
PRAG 30.82 0.56 (0.47x) 672 (1x)
DyPRAG ($p=2$) 32.66 0.56+0.13+0.060 (0.625x) 7.71 (0.011x)
DyPRAG ($p=4$) 33.26 0.56+0.13+0.062 (0.627x) 15.42 (0.023x)
DyPRAG ($p=16$) 32.08 0.56+0.13+0.055 (0.621x) 61.70 (0.092x)
DyPRAG ($p=32$) 31.94 0.56+0.13+0.060 (0.625x) 123.39 (0.184x)
@@
By combining in-context injection with parameter injection, DyPRAG-Combine achieves the best performance across all models, outperforming all baselines. For instance, DyPRAG-Combine outperforms PRAG-Combine by 1.86% on LLaMA3.2-1B, 0.46% on Qwen2.5-1.5B and 1.08% on LLaMA3-8B on average. These results demonstrate that the dynamic parameters introduced by our approach effectively intergrade with contextual knowledge, enabling the two information sources to complement each other.
@@ 4.4 Out-Of-Distribution (OOD) Performance
To further demonstrate the generalization ability of the DyPRAG method, we evaluate it in an out-of-distribution (OOD) scenario. Table 3 presents the OOD performance on the StrategyQA (SQA)[12] and IIRC[9] datasets. Compared to the standard RAG model, the vanilla model performs poorly due to the lack of relevant knowledge. In contrast, our DyPRAG successfully injects parametric knowledge, achieving comparable performance in SQA. Notably, DyPRAG-Combine which incorporates document-related parametric knowledge with ground-truth passages can answer these questions more correctly, achieving the best performance across all scenarios. For example, DyPRAG-Combine improves performance on SQA by 22.66% using LLaMA3.2-1B and on IIRC by 13.63% using LLaMA3-8B. These experimental results highlight the significant generalization ability of DyPRAG, surpassing that of PRAG, which is unable to handle this OOD scenario without additional offline training. Detailed implementation is provided in Appendix B.1 and analysis in Appendix D.
4.5 Ablation Study
Effect of Intermediate Dimension $p$.
When contextual knowledge and parametric knowledge conflict, LLMs may struggle to determine which source of information is more reliable [42; 51]. A key distinction between parametric RAG and standard RAG lies in whether external documents are used as contextual knowledge or transforming into parametric knowledge. Additionally, we aim to explore why combining parameter injection with in-context injection results in superior performance.